OpenStudy (ttop0816):

please help! If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. What is the numerator of the fraction? 2 9 16

OpenStudy (solomonzelman):

\[\frac{y}{x+3}=\frac{3}{4}\]\[\frac{y}{x-7}=1\] you can conclude.\[y=x-7\]So,\[\frac{x-7}{x+3}=\frac{3}{4}\]cross multiplication.\[4(x-7)=3(x+3)\]\[4x-28=3x+9\]SOLVE FOR x then tell me what x+3 is going to be equal to.

OpenStudy (ttop0816):

x=37

OpenStudy (ttop0816):

then do i just plug in 37 in to x+3??

OpenStudy (solomonzelman):

to find the denominator yes.

OpenStudy (ttop0816):

but im trying to find the numerator;

OpenStudy (solomonzelman):

oh, then plug x=37 into x-7

OpenStudy (solomonzelman):

You might want to go over my work...

OpenStudy (solomonzelman):

then idk.